# Multiview Random Vector Functional Link Network for Predicting DNA-Binding Proteins

### v1

@@ A.3.1 Sensitivity of $\theta$ and $\rho$ combinations
Table 9: Performance comparison of the proposed MvRVFL along with the baseline models based on classification Acc. for Corel5k datasets.
Dataset SVM2K MvTSVM ELM-Vw-$A$ ELM-Vw-$B$ RVFL-Vw-$A$ RVFL-Vw-$B$ MVLDM MvRVFL-1 MvRVFL-2
(Farquhar et al., 2005) (Xie & Sun, 2015) (Huang et al., 2006) (Huang et al., 2006) (Pao et al., 1994) (Pao et al., 1994) (Hu et al., 2024)
$(C_{1})$ $(C_{1},C_{2},D)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C_{1},\nu_{1},\nu_{2},\sigma)$ $(C_{1},C_{2},\rho,N)$ $(C_{1},C_{2},\rho,N)$
1000 $81.67$ $50$ $83.33$ $78.33$ $83.33$ $76.67$ $73.33$ $83.33$ $85$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,43)$ $(0.0001,183)$ $(0.00001,3)$ $(0.01,0.001,0.001,4)$ $(0.001,0.01,0.00001,23)$ $(0.0001,100000,0.00001,203)$
10000 $71.67$ $48.33$ $56.67$ $63.33$ $71.67$ $71.67$ $55$ $71.67$ $70$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(1000,23)$ $(1000,23)$ $(0.01,3)$ $(0.00001,63)$ $(0.01,0.01,0.001,4)$ $(0.01,1000,0.01,23)$ $(0.001,100,0.0001,143)$
100000 $71.67$ $55$ $63.33$ $73.33$ $78.33$ $75$ $61.67$ $65$ $71.67$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(100000,183)$ $(0.00001,123)$ $(0.001,183)$ $(0.00001,123)$ $(100,1000,100,4)$ $(0.1,10000,0.01,143)$ $(1,100000,0.001,143)$
101000 $66.67$ $58.33$ $75$ $73.33$ $68.33$ $76.67$ $55$ $81.67$ $73.33$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.001,123)$ $(0.00001,143)$ $(0.001,3)$ $(0.00001,143)$ $(100,10,100,4)$ $(0.001,1000,0.00001,103)$ $(1,100000,0.001,23)$
102000 $81.67$ $51.67$ $81.67$ $75$ $81.67$ $71.67$ $83.33$ $78.33$ $76.67$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.01,63)$ $(0.00001,143)$ $(0.01,3)$ $(100000,3)$ $(0.001,0.001,0.0001,4)$ $(0.001,10,0.00001,23)$ $(0.001,10,0.00001,43)$
103000 $73.33$ $45$ $80$ $75$ $81.67$ $73.33$ $65$ $81.67$ $80$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.0001,43)$ $(0.001,163)$ $(0.00001,43)$ $(0.0001,0.1,0.001,4)$ $(0.01,1000,0.001,23)$ $(0.1,10000,0.01,23)$
104000 $71.67$ $46.67$ $55$ $65$ $76.67$ $76.67$ $46.67$ $71.67$ $53.33$
$(32)$ $(0.00001,0.00001,0.00001)$ $(10,123)$ $(10000,23)$ $(0.0001,203)$ $(0.00001,83)$ $(0.0001,0.0001,0.00001,4)$ $(100000,0.00001,0.001,83)$ $(1000,100000,0.00001,23)$
108000 $80$ $48.33$ $75.65$ $76.67$ $78.33$ $76.67$ $78.33$ $83.33$ $80$
$(8)$ $(0.00001,0.00001,0.00001)$ $(0.001,63)$ $(0.00001,83)$ $(0.1,3)$ $(0.00001,3)$ $(100,1000,10000,4)$ $(0.001,0.00001,0.00001,163)$ $(0.0001,10,0.00001,3)$
109000 $65$ $43.33$ $68.33$ $73.33$ $68.33$ $75$ $73.33$ $75$ $58.33$
$(1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(100,63)$ $(0.0001,203)$ $(0.0001,43)$ $(0.001,0.001,0.0001,0.25)$ $(0.01,10000,0.00001,203)$ $(0.01,100000,0.0001,83)$
113000 $66.67$ $45$ $68.33$ $63.33$ $70$ $63.33$ $66.67$ $70$ $68.33$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,123)$ $(1000,203)$ $(0.001,123)$ $(0.0001,23)$ $(0.0001,0.001,0.001,0.25)$ $(0.01,0.00001,0.001,3)$ $(0.01,0.001,0.00001,3)$
118000 $58.33$ $48.33$ $68.33$ $61.67$ $68.33$ $56.67$ $58.33$ $71.67$ $53.33$
$(8)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(100000,43)$ $(0.001,103)$ $(1000,43)$ $(0.01,0.1,0.01,4)$ $(100000,100000,1,63)$ $(1,100000,0.0001,143)$
119000 $71.33$ $53.33$ $81.67$ $83.33$ $78.33$ $80$ $70$ $78.33$ $71.67$
$(16)$ $(0.00001,0.00001,0.00001)$ $(0.001,143)$ $(0.00001,123)$ $(0.01,63)$ $(0.00001,83)$ $(0.001,0.0001,0.01,0.25)$ $(1,100000,0.1,103)$ $(0.1,100000,0.01,123)$
12000 $60$ $48.33$ $63.33$ $58.33$ $63.33$ $51.67$ $61.67$ $61.67$ $56.67$
$(32)$ $(0.00001,0.00001,0.00001)$ $(0.0001,83)$ $(0.00001,63)$ $(0.0001,83)$ $(0.00001,43)$ $(0.1,0.1,0.1,2)$ $(0.0001,1000,0.00001,143)$ $(0.00001,100,0.00001,203)$
120000 $73.33$ $48.33$ $78.33$ $80$ $78.33$ $85$ $66.67$ $75$ $88.33$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.01,203)$ $(0.00001,203)$ $(0.01,203)$ $(0.00001,103)$ $(0.001,0.0001,0.001,4)$ $(0.0001,10,100,203)$ $(0.1,1000,0.01,23)$
121000 $63.33$ $55$ $68.33$ $75$ $71.67$ $71.67$ $71.67$ $73.33$ $83.33$
$(2)$ $(0.00001,0.00001,0.00001)$ $(0.0001,163)$ $(0.00001,103)$ $(0.001,143)$ $(0.00001,103)$ $(0.001,0.1,0.01,2)$ $(0.01,10000,0.00001,83)$ $(0.001,1000,0.00001,63)$
122000 $78.33$ $55$ $70$ $76.67$ $70$ $71.67$ $61.67$ $80$ $78.33$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.01,163)$ $(100,23)$ $(0.01,163)$ $(0.00001,123)$ $(0.01,0.001,0.01,0.25)$ $(0.0001,100,0.0001,163)$ $(0.01,10000,0.00001,83)$
13000 $85$ $50$ $83.33$ $91.67$ $85$ $90$ $78.33$ $85$ $86.67$
$(32)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(0.00001,103)$ $(0.001,183)$ $(0.00001,43)$ $(0.01,0.001,0.0001,2)$ $(0.1,1000,0.00001,3)$ $(0.001,100,0.0001,163)$
130000 $58.96$ $48.33$ $58.33$ $65$ $58.33$ $61.67$ $60$ $61.67$ $66.67$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(100,23)$ $(0.001,103)$ $(0.00001,23)$ $(0.001,0.001,0.001,4)$ $(0.00001,1000,0.0001,103)$ $(0.01,100000,0.0001,163)$
131000 $78.33$ $63.33$ $73.33$ $76.67$ $83.33$ $76.67$ $81.67$ $80$ $78.33$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.00001,83)$ $(0.01,23)$ $(0.00001,183)$ $(0.01,0.001,0.01,4)$ $(0.001,1000,0.00001,123)$ $(10,100000,0.1,203)$
140000 $71.67$ $40$ $58.33$ $76.67$ $56.67$ $78.33$ $45$ $78.33$ $66.67$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.1,183)$ $(0.00001,143)$ $(0.1,3)$ $(100,3)$ $(0.01,0.001,0.0001,4)$ $(1000,0.001,10,3)$ $(10000,100,0.0001,3)$
142000 $86.67$ $51.67$ $85$ $68.33$ $91.67$ $88.33$ $78.33$ $91.67$ $90$
$(0.25)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(10000,23)$ $(0.1,3)$ $(1,3)$ $(0.01,0.001,0.0001,4)$ $(0.01,0.0001,0.001,3)$ $(0.1,10,0.0001,3)$
143000 $66.67$ $51.67$ $63.33$ $65$ $63.33$ $65$ $71.67$ $60$ $60$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.00001,203)$ $(0.001,163)$ $(0.00001,203)$ $(0.01,0.1,0.1,0.25)$ $(0.001,0.1,0.0001,203)$ $(10,0.001,0.001,3)$
144000 $73.33$ $48.33$ $68.33$ $71.67$ $68.33$ $70$ $66.67$ $76.67$ $73.33$
$(16)$ $(0.00001,0.00001,0.00001)$ $(0.01,63)$ $(0.00001,203)$ $(0.01,63)$ $(0.00001,203)$ $(0.1,0.1,0.1,0.25)$ $(0.001,1000,0.00001,143)$ $(0.001,1000,0.00001,43)$
147000 $58.33$ $55$ $63.33$ $73.33$ $63.33$ $73.33$ $70$ $71.67$ $66.67$
$(0.125)$ $(0.00001,0.00001,0.00001)$ $(0.01,203)$ $(0.00001,123)$ $(0.01,203)$ $(0.00001,123)$ $(0.001,0.001,0.1,4)$ $(0.1,1,0.00001,203)$ $(0.01,100000,0.00001,103)$
148000 $86.67$ $50$ $90$ $80$ $88.33$ $86.67$ $83.33$ $88.33$ $83.33$
$(2)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(0.00001,83)$ $(0.001,23)$ $(0.00001,63)$ $(0.001,0.0001,0.1,4)$ $(0.001,100,0.00001,123)$ $(0.001,1000,0.00001,83)$
152000 $56.67$ $48.33$ $65$ $66.67$ $66.67$ $45$ $51.67$ $66.67$ $66.67$
$(0.25)$ $(0.00001,0.00001,0.00001)$ $(0.0001,143)$ $(0.1,23)$ $(0.00001,103)$ $(1,103)$ $(0.1,0.1,0.001,0.25)$ $(0.00001,1000,0.00001,163)$ $(0.001,100000,0.0001,3)$
153000 $79.33$ $55$ $76.67$ $73.33$ $76.67$ $68.33$ $80$ $81.67$ $80$
$(0.5)$ $(0.00001,0.00001,0.00001)$ $(0.001,203)$ $(0.00001,43)$ $(0.001,203)$ $(0.00001,43)$ $(0.01,10,100,4)$ $(0.01,0.00001,0.001,83)$ $(0.01,100,0.001,103)$
161000 $86.67$ $46.67$ $91.67$ $91.67$ $88.33$ $90$ $93.33$ $93.33$ $91.67$
$(0.5)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(0.00001,143)$ $(0.1,23)$ $(0.00001,43)$ $(0.001,0.0001,0.00001,2)$ $(0.1,0.01,0.00001,23)$ $(0.01,1000,0.001,143)$
163000 $80$ $41.67$ $80$ $78.33$ $80$ $85$ $71.67$ $90$ $90$
$(4)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(1,23)$ $(0.001,83)$ $(0.00001,3)$ $(0.001,0.0001,0.00001,0.5)$ $(0.01,1000,0.00001,43)$ $(0.1,100000,0.00001,3)$
17000 $83.33$ $56.67$ $86.67$ $90$ $86.67$ $86.67$ $83.33$ $91.67$ $83.33$
$(2)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.00001,143)$ $(0.001,163)$ $(0.00001,203)$ $(0.001,0.00001,0.1,4)$ $(0.01,1000,0.001,183)$ $(0.01,10000,0.00001,63)$
171000 $88.33$ $48.33$ $76.67$ $68.33$ $76.67$ $75$ $63.33$ $80$ $70$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,123)$ $(0.00001,83)$ $(0.001,123)$ $(0.00001,3)$ $(0.1,0.001,0.1,0.25)$ $(0.0001,100,0.0001,103)$ $(0.001,100,0.0001,103)$
173000 $85$ $61.67$ $86.67$ $71.67$ $83.33$ $78.33$ $80$ $86.67$ $80$
$(0.125)$ $(0.00001,0.00001,0.00001)$ $(0.001,203)$ $(1000,23)$ $(0.001,63)$ $(1000,3)$ $(0.0001,0.01,0.1,0.25)$ $(0.001,10000,0.0001,183)$ $(0.001,10000,0.00001,123)$
174000 $81.67$ $45$ $81.67$ $85$ $83.33$ $88.33$ $88.33$ $93.33$ $83.33$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.01,103)$ $(0.001,43)$ $(0.01,103)$ $(0.00001,23)$ $(0.0001,0.01,0.0001,4)$ $(1,100000,0.001,23)$ $(0.1,10000,0.00001,203)$
182000 $76.67$ $46.67$ $78.33$ $76.67$ $80$ $76.67$ $70$ $81.67$ $75$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(0.00001,123)$ $(0.001,183)$ $(0.00001,123)$ $(0.00001,0.0001,0.0001,0.25)$ $(0.01,0.00001,0.001,3)$ $(0.01,100,0.00001,3)$
183000 $70$ $48.33$ $80$ $68.33$ $80$ $78.33$ $68.33$ $80$ $78.33$
[…]

### v2

@@ Appendix B Random Vector Functional Link (RVFL) Network
Table S.III: Performance comparison of the proposed MvRVFL along with the baseline models based on classification Acc. for Corel5k datasets.
Dataset SVM2K [18] MvTSVM [70] ELM-Vw-AA [31] ELM-Vw-$B$ [31] RVFL-Vw-AA [46] RVFL-Vw-$B$ [46] MVLDM [29] CDC [34] MCMC [5] MvRVFL-1 MvRVFL-2
$(C_{1})$ $(C_{1},C_{2},D)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C,N)$ $(C_{1},\nu_{1},\nu_{2},\sigma)$ $(\alpha,\beta,m)$ $(\lambda_{1},\lambda_{2},\lambda_{3},\zeta)$ $(C_{1},C_{2},\rho,N)$ $(C_{1},C_{2},\rho,N)$
1000 $81.67$ $(50)$ $(83.33)$ $(78.33)$ $(83.33)$ $(76.67)$ $(73.33)$ $82.17$ $82$ $(83.33)$ $(85)$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.0001,183)$ $(0.00001,43)$ $(0.0001,183)$ $(0.00001,3)$ $(0.25,0.25,0.25,4)$ $(1,10,100)$ $(0.01,0.1,0.000001,0.000001)$ $(0.001,0.01,0.00001,23)$ $(0.0001,100000,0.00001,203)$
10000 $71.67$ $(48.33)$ $(56.67)$ $(63.33)$ $(71.67)$ $(71.67)$ $(55)$ $70.87$ $69.42$ $(71.67)$ $(70)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(1000,23)$ $(1000,23)$ $(0.01,3)$ $(0.00001,63)$ $(0.25,0.25,0.25,4)$ $(1000,0.0001,50)$ $(0.1,0.01,0.00001,0.000001)$ $(0.01,1000,0.01,23)$ $(0.001,100,0.0001,143)$
100000 $71.67$ $(55)$ $(63.33)$ $(73.33)$ $(78.33)$ $(75)$ $(61.67)$ $62.45$ $63.25$ $(65)$ $(71.67)$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(100000,183)$ $(0.00001,123)$ $(0.001,183)$ $(0.00001,123)$ $(0.25,0.25,0.25,4)$ $(1000,10,100)$ $(0.1,1,0.0001,0.0001)$ $(0.1,10000,0.01,143)$ $(1,100000,0.001,143)$
101000 $66.67$ $(58.33)$ $(75)$ $(73.33)$ $(68.33)$ $(76.67)$ $(55)$ $72.49$ $71.47$ $(81.67)$ $(73.33)$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.001,123)$ $(0.00001,143)$ $(0.001,3)$ $(0.00001,143)$ $(0.25,0.25,0.25,4)$ $(10000,10000,10)$ $(0.01,1,0.000001,0.00001)$ $(0.001,1000,0.00001,103)$ $(1,100000,0.001,23)$
102000 $81.67$ $(51.67)$ $(81.67)$ $(75)$ $(81.67)$ $(71.67)$ $(83.33)$ $75.04$ $73.5$ $(78.33)$ $(76.67)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.01,63)$ $(0.00001,143)$ $(0.01,3)$ $(100000,3)$ $(0.25,0.25,0.25,4)$ $(1,10000,70)$ $(2.5,0.01,0.000001,0.00001)$ $(0.001,10,0.00001,23)$ $(0.001,10,0.00001,43)$
103000 $73.33$ $(45)$ $(80)$ $(75)$ $(81.67)$ $(73.33)$ $(65)$ $80.86$ $79$ $(81.67)$ $(80)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.0001,43)$ $(0.001,163)$ $(0.00001,43)$ $(0.25,0.25,0.25,4)$ $(1,0.0001,100)$ $(5,0.1,0.001,0.0001)$ $(0.01,1000,0.001,23)$ $(0.1,10000,0.01,23)$
104000 $71.67$ $(46.67)$ $(55)$ $(65)$ $(76.67)$ $(76.67)$ $(46.67)$ $69.58$ $68.5$ $(71.67)$ $(53.33)$
$(32)$ $(0.00001,0.00001,0.00001)$ $(10,123)$ $(10000,23)$ $(0.0001,203)$ $(0.00001,83)$ $(0.25,0.25,0.25,4)$ $(0.0001,0.0001,100)$ $(0.1,1,0.00001,0.001)$ $(100000,0.00001,0.001,83)$ $(1000,100000,0.00001,23)$
108000 $80$ $(48.33)$ $(75.65)$ $(76.67)$ $(78.33)$ $(76.67)$ $(78.33)$ $79.75$ $78.5$ $(83.33)$ $(80)$
$(8)$ $(0.00001,0.00001,0.00001)$ $(0.001,63)$ $(0.00001,83)$ $(0.1,3)$ $(0.00001,3)$ $(0.25,0.25,0.25,4)$ $(0.0001,0.0001,10)$ $(1,0.1,0.0001,0.000001)$ $(0.001,0.00001,0.00001,163)$ $(0.0001,10,0.00001,3)$
109000 $65$ $(43.33)$ $(68.33)$ $(73.33)$ $(68.33)$ $(75)$ $(73.33)$ $57.28$ $55.5$ $(75)$ $(58.33)$
$(1)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(100,63)$ $(0.0001,203)$ $(0.0001,43)$ $(0.25,0.25,0.25,0.25)$ $(0.0001,0.0001,30)$ $(10,1,0.001,0.00001)$ $(0.01,10000,0.00001,203)$ $(0.01,100000,0.0001,83)$
113000 $66.67$ $(45)$ $(68.33)$ $(63.33)$ $(70)$ $(63.33)$ $(66.67)$ $66.63$ $65.5$ $(70)$ $(68.33)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,123)$ $(1000,203)$ $(0.001,123)$ $(0.0001,23)$ $(0.25,0.25,0.25,0.25)$ $(0.0001,1,10)$ $(0.01,5,0.00001,0.0001)$ $(0.01,0.00001,0.001,3)$ $(0.01,0.001,0.00001,3)$
118000 $58.33$ $(48.33)$ $(68.33)$ $(61.67)$ $(68.33)$ $(56.67)$ $(58.33)$ $70.18$ $69.5$ $(71.67)$ $(53.33)$
$(8)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(100000,43)$ $(0.001,103)$ $(1000,43)$ $(0.25,0.25,0.25,4)$ $(0.0001,1,30)$ $(2.5,2.5,0.0001,0.001)$ $(100000,100000,1,63)$ $(1,100000,0.0001,143)$
119000 $71.33$ $(53.33)$ $(81.67)$ $(83.33)$ $(78.33)$ $(80)$ $(70)$ $70.78$ $69.87$ $(78.33)$ $(71.67)$
$(16)$ $(0.00001,0.00001,0.00001)$ $(0.001,143)$ $(0.00001,123)$ $(0.01,63)$ $(0.00001,83)$ $(0.25,0.25,0.25,0.25)$ $(0.0001,10000,10)$ $(5,2.5,0.000001,0.00001)$ $(1,100000,0.1,103)$ $(0.1,100000,0.01,123)$
12000 $60$ $(48.33)$ $(63.33)$ $(58.33)$ $(63.33)$ $(51.67)$ $(61.67)$ $60.55$ $59.23$ $(61.67)$ $(56.67)$
$(32)$ $(0.00001,0.00001,0.00001)$ $(0.0001,83)$ $(0.00001,63)$ $(0.0001,83)$ $(0.00001,43)$ $(0.25,0.25,0.25,2)$ $(0.0001,10000,30)$ $(0.1,10,0.001,0.0001)$ $(0.0001,1000,0.00001,143)$ $(0.00001,100,0.00001,203)$
120000 $73.33$ $(48.33)$ $(78.33)$ $(80)$ $(78.33)$ $(85)$ $(66.67)$ $74.83$ $74$ $(75)$ $(88.33)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.01,203)$ $(0.00001,203)$ $(0.01,203)$ $(0.00001,103)$ $(0.25,0.25,0.25,4)$ $(1,0.0001,10)$ $(1,0.01,0.00001,0.001)$ $(0.0001,10,100,203)$ $(0.1,1000,0.01,23)$
121000 $63.33$ $(55)$ $(68.33)$ $(75)$ $(71.67)$ $(71.67)$ $(71.67)$ $72.35$ $71.5$ $(73.33)$ $(83.33)$
$(2)$ $(0.00001,0.00001,0.00001)$ $(0.0001,163)$ $(0.00001,103)$ $(0.001,143)$ $(0.00001,103)$ $(0.25,0.25,0.25,2)$ $(1,0.0001,30)$ $(10,0.1,0.0001,0.000001)$ $(0.01,10000,0.00001,83)$ $(0.001,1000,0.00001,63)$
122000 $78.33$ $(55)$ $(70)$ $(76.67)$ $(70)$ $(71.67)$ $(61.67)$ $77.85$ $75$ $(80)$ $(78.33)$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.01,163)$ $(100,23)$ $(0.01,163)$ $(0.00001,123)$ $(0.25,0.25,0.25,0.25)$ $(1,1,10)$ $(0.01,0.01,0.001,0.00001)$ $(0.0001,100,0.0001,163)$ $(0.01,10000,0.00001,83)$
13000 $85$ $(50)$ $(83.33)$ $(91.67)$ $(85)$ $(90)$ $(78.33)$ $86.95$ $86$ $(85)$ $(86.67)$
$(32)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(0.00001,103)$ $(0.001,183)$ $(0.00001,43)$ $(0.25,0.25,0.25,2)$ $(1,1,30)$ $(2.5,5,0.00001,0.0001)$ $(0.1,1000,0.00001,3)$ $(0.001,100,0.0001,163)$
130000 $58.96$ $(48.33)$ $(58.33)$ $(65)$ $(58.33)$ $(61.67)$ $(60)$ $60.51$ $65.5$ $(61.67)$ $(66.67)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(100,23)$ $(0.001,103)$ $(0.00001,23)$ $(0.25,0.25,0.25,4)$ $(1,10,10)$ $(5,0.01,0.0001,0.001)$ $(0.00001,1000,0.0001,103)$ $(0.01,100000,0.0001,163)$
131000 $78.33$ $(63.33)$ $(73.33)$ $(76.67)$ $(83.33)$ $(76.67)$ $(81.67)$ $75.03$ $77$ $(80)$ $(78.33)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.00001,83)$ $(0.01,23)$ $(0.00001,183)$ $(0.25,0.25,0.25,4)$ $(1,10,30)$ $(0.1,2.5,0.000001,0.00001)$ $(0.001,1000,0.00001,123)$ $(10,100000,0.1,203)$
140000 $71.67$ $(40)$ $(58.33)$ $(76.67)$ $(56.67)$ $(78.33)$ $(45)$ $66.44$ $65.5$ $(78.33)$ $(66.67)$
$(0.0625)$ $(0.00001,0.00001,0.00001)$ $(0.1,183)$ $(0.00001,143)$ $(0.1,3)$ $(100,3)$ $(0.25,0.25,0.25,4)$ $(1,1000,10)$ $(1,1,0.001,0.0001)$ $(1000,0.001,10,3)$ $(10000,100,0.0001,3)$
142000 $86.67$ $(51.67)$ $(85)$ $(68.33)$ $(91.67)$ $(88.33)$ $(78.33)$ $90.44$ $89.5$ $(91.67)$ $(90)$
$(0.25)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(10000,23)$ $(0.1,3)$ $(1,3)$ $(0.25,0.25,0.25,4)$ $(1,1000,30)$ $(10,2.5,0.00001,0.001)$ $(0.01,0.0001,0.001,3)$ $(0.1,10,0.0001,3)$
143000 $66.67$ $(51.67)$ $(63.33)$ $(65)$ $(63.33)$ $(65)$ $(71.67)$ $59.7$ $57$ $(60)$ $(60)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.00001,203)$ $(0.001,163)$ $(0.00001,203)$ $(0.25,0.25,0.25,0.25)$ $(1,10000,10)$ $(0.01,0.1,0.0001,0.000001)$ $(0.001,0.1,0.0001,203)$ $(10,0.001,0.001,3)$
144000 $73.33$ $(48.33)$ $(68.33)$ $(71.67)$ $(68.33)$ $(70)$ $(66.67)$ $72.5$ $72.5$ $(76.67)$ $(73.33)$
$(16)$ $(0.00001,0.00001,0.00001)$ $(0.01,63)$ $(0.00001,203)$ $(0.01,63)$ $(0.00001,203)$ $(0.25,0.25,0.25,0.25)$ $(1,10000,30)$ $(2.5,10,0.001,0.00001)$ $(0.001,1000,0.00001,143)$ $(0.001,1000,0.00001,43)$
147000 $58.33$ $(55)$ $(63.33)$ $(73.33)$ $(63.33)$ $(73.33)$ $(70)$ $66.84$ $64.5$ $(71.67)$ $(66.67)$
$(0.125)$ $(0.00001,0.00001,0.00001)$ $(0.01,203)$ $(0.00001,123)$ $(0.01,203)$ $(0.00001,123)$ $(0.25,0.25,0.25,4)$ $(10,0.0001,10)$ $(5,5,0.00001,0.0001)$ $(0.1,1,0.00001,203)$ $(0.01,100000,0.00001,103)$
148000 $86.67$ $(50)$ $(90)$ $(80)$ $(88.33)$ $(86.67)$ $(83.33)$ $81.83$ $85.5$ $(88.33)$ $(83.33)$
$(2)$ $(0.00001,0.00001,0.00001)$ $(0.001,103)$ $(0.00001,83)$ $(0.001,23)$ $(0.00001,63)$ $(0.25,0.25,0.25,4)$ $(10,0.0001,30)$ $(0.1,0.01,0.0001,0.001)$ $(0.001,100,0.00001,123)$ $(0.001,1000,0.00001,83)$
152000 $56.67$ $(48.33)$ $(65)$ $(66.67)$ $(66.67)$ $(45)$ $(51.67)$ $65.37$ $63$ $(66.67)$ $(66.67)$
$(0.25)$ $(0.00001,0.00001,0.00001)$ $(0.0001,143)$ $(0.1,23)$ $(0.00001,103)$ $(1,103)$ $(0.25,0.25,0.25,0.25)$ $(10,1,10)$ $(1,2.5,0.000001,0.00001)$ $(0.00001,1000,0.00001,163)$ $(0.001,100000,0.0001,3)$
153000 $79.33$ $(55)$ $(76.67)$ $(73.33)$ $(76.67)$ $(68.33)$ $(80)$ $79.58$ $77$ $(81.67)$ $(80)$
$(0.5)$ $(0.00001,0.00001,0.00001)$ $(0.001,203)$ $(0.00001,43)$ $(0.001,203)$ $(0.00001,43)$ $(0.25,0.25,0.25,4)$ $(10,1,30)$ $(10,0.01,0.001,0.0001)$ $(0.01,0.00001,0.001,83)$ $(0.01,100,0.001,103)$
161000 $86.67$ $(46.67)$ $(91.67)$ $(91.67)$ $(88.33)$ $(90)$ $(93.33)$ $90.91$ $91$ $(93.33)$ $(91.67)$
$(0.5)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(0.00001,143)$ $(0.1,23)$ $(0.00001,43)$ $(0.25,0.25,0.25,2)$ $(10,10,30)$ $(0.01,1,0.00001,0.001)$ $(0.1,0.01,0.00001,23)$ $(0.01,1000,0.001,143)$
163000 $80$ $(41.67)$ $(80)$ $(78.33)$ $(80)$ $(85)$ $(71.67)$ $89.41$ $86$ $(90)$ $(90)$
$(4)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(1,23)$ $(0.001,83)$ $(0.00001,3)$ $(0.25,0.25,0.25,0.5)$ $(1,1,10)$ $(2.5,0.1,0.0001,0.000001)$ $(0.01,1000,0.00001,43)$ $(0.1,100000,0.00001,3)$
17000 $83.33$ $(56.67)$ $(86.67)$ $(90)$ $(86.67)$ $(86.67)$ $(83.33)$ $82.36$ $82.5$ $(91.67)$ $(83.33)$
$(2)$ $(0.00001,0.00001,0.00001)$ $(0.001,163)$ $(0.00001,143)$ $(0.001,163)$ $(0.00001,203)$ $(0.25,0.25,0.25,4)$ $(1,1000,100)$ $(5,1,0.001,0.00001)$ $(0.01,1000,0.001,183)$ $(0.01,10000,0.00001,63)$
171000 $88.33$ $(48.33)$ $(76.67)$ $(68.33)$ $(76.67)$ $(75)$ $(63.33)$ $69.63$ $68.47$ $(80)$ $(70)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,123)$ $(0.00001,83)$ $(0.001,123)$ $(0.00001,3)$ $(0.25,0.25,0.25,0.25)$ $(0.0001,0.0001,10)$ $(0.1,5,0.00001,0.0001)$ $(0.0001,100,0.0001,103)$ $(0.001,100,0.0001,103)$
173000 $85$ $(61.67)$ $(86.67)$ $(71.67)$ $(83.33)$ $(78.33)$ $(80)$ $79.61$ $78.5$ $(86.67)$ $(80)$
$(0.125)$ $(0.00001,0.00001,0.00001)$ $(0.001,203)$ $(1000,23)$ $(0.001,63)$ $(1000,3)$ $(0.25,0.25,0.25,0.25)$ $(10,10,10)$ $(1,10,0.0001,0.001)$ $(0.001,10000,0.0001,183)$ $(0.001,10000,0.00001,123)$
174000 $81.67$ $(45)$ $(81.67)$ $(85)$ $(83.33)$ $(88.33)$ $(88.33)$ $82.15$ $80.5$ $(93.33)$ $(83.33)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.01,103)$ $(0.001,43)$ $(0.01,103)$ $(0.00001,23)$ $(0.25,0.25,0.25,4)$ $(1,10000,10)$ $(10,5,0.000001,0.00001)$ $(1,100000,0.001,23)$ $(0.1,10000,0.00001,203)$
182000 $76.67$ $(46.67)$ $(78.33)$ $(76.67)$ $(80)$ $(76.67)$ $(70)$ $74.15$ $75$ $(81.67)$ $(75)$
$(0.03125)$ $(0.00001,0.00001,0.00001)$ $(0.001,183)$ $(0.00001,123)$ $(0.001,183)$ $(0.00001,123)$ $(0.25,0.25,0.25,0.25)$ $(1,0.0001,10)$ $(0.01,2.5,0.001,0.0001)$ $(0.01,0.00001,0.001,3)$ $(0.01,100,0.00001,3)$
183000 $70$ $(48.33)$ $(80)$ $(68.33)$ $(80)$ $(78.33)$ $(68.33)$ $77.63$ $77$ $(80)$ $(78.33)$
$(4)$ $(0.00001,0.00001,0.00001)$ $(0.0001,203)$ $(1000,43)$ $(0.0001,203)$ $(0.00001,43)$ $(0.25,0.25,0.25,4)$ $(0.0001,1000,10)$ $(2.5,0.01,0.00001,0.001)$ $(0.01,0.01,0.001,23)$ $(0.001,100,0.00001,143)$
[…]